OpenStudy (anonymous):

Solve the system using an augmented matrix. Please show all your steps! Thanks! (: -x -3y = -17 2x -6y = 26

OpenStudy (anonymous):

-1 -3 l -17 2 -6 l 26 Row1 x2 + row 2 = -1 -3 l -17 0 -12 l -8 -12y = -8 y = -(2/3)

OpenStudy (anonymous):

Just kidding y = 2/3